Find the minimum or maximum value of the quadratic function g(x)=2x^2+6x+3
Answer by scott8148(6628)

the minimum (or maximum) lies on the axis of symmetry

the general eqn for the axis is x=-b/(2a)

in this case, x=-6/(2*2) __ x=-3/2

so the MINIMUM value (because a is positive) is 2(-3/2)^2+6(-3/2)+3 or -3/2